The circle is tangent to the line 4x+3y=4 at the point (4,-4) and center is on the line x-y=7
Answer by ikleyn(52908)

.

A good style (and the REQUIREMENT of this forum) is to ask a question in each post you submit to this forum.

So, what is your question (in case if you have it .. . ) ?


----------------
Comment from student : The question is, what is the equation of the circle. And how will I get the center to the given equation x-y=7.
----------------


My response :

1.  Find equation of the straight line perpendicular to the given line 4x + 3y = 4  and passing through the point (4,-4).



2.  Find the intersection of that perpendicular line with the other given line  x - y = 7.


     For it, you need to solve a system of 2 linear equations in 2 unknown.



3.  The intersection point will be the center of the circle the problems asks for.



4.  The radius of the circle will be the distance from the center to the given point (4,-4).


Following these instructions, you can solve the problem on your own.